OpenStudy (anonymous):

What is the solution to the equation? Medal for help :) -18 + 4 + x = 9 - 19 x = ? <<

OpenStudy (anonymous):

-14 + x = -28 +14 +14 x = -14

OpenStudy (popitree):

its -14+x = -10

OpenStudy (popitree):

x=4

OpenStudy (anonymous):

thank u @pondamo

OpenStudy (anonymous):

@annafornia wait, i made a mistake, sorry @popitree is correct

OpenStudy (anonymous):

how did u get that @popitree

OpenStudy (anonymous):

-14 + x = -10 +14 +14 x = 4

OpenStudy (anonymous):

thank u! @popitree & @pondamo haha :)))

OpenStudy (anonymous):

@annafornia yw, sorry for the mistake xD